OpenStudy (anonymous):

evaluate the limit (lim x-->pi/2 ) sec x

OpenStudy (anonymous):

as \(x\to \frac{\pi}{2}\) you have \(\cos(x)\to 0\)

OpenStudy (anonymous):

since \(\frac{1}{0}\) is undefined, there is no limit